OpenStudy (samsan9):

solve for x. write your answer to three decimal places. (1000)(1.05)^3x=3000

OpenStudy (anonymous):

divide both sides by 1000, what do you have?

OpenStudy (samsan9):

(1.05)^3x=3

OpenStudy (anonymous):

now, apply the rule a^b =c , then log(base a) of c =b what do you have?

OpenStudy (samsan9):

1.05^3=3x?

OpenStudy (anonymous):

not that, log

OpenStudy (anonymous):